OpenStudy (anonymous):

How long should Bill's boss plan for him to cover an 8.9 mile route if Bill averages 2.1 mph? Round to the nearest tenth if necessary.

OpenStudy (mathstudent55):

average speed = distance/time v = d/t Solving for t, you get: t = d/v time = distance/average speed

OpenStudy (anonymous):

so how do you do it ?

OpenStudy (mathstudent55):

Divide the distance by the average speed.